the weight of meteorite a is 3 times the weight of meterorite b . If the sum of their weights is 76 tons,find the weight of each.
Answer by amruta(4)

let the weight of m. a be x
hence, weight of m. b will be 3x
i.e. x+3x=76 tons
4x=76
x=19 tons.
meteorite a = 19 tons
" b = 19*3= 57 tons.